Help please
6.7.4 Current Density and Drift Speed (a) A group of charges, each with charge q, moves with velocity v. The number of particles per unit volume is n. What is the current density J of these charges, in magnitude and direction? Make sure that your answer has units of A/m2 (b) We want to calculate how long it takes an electron to get from a car battery to the starter motor after the ignition switch is turned. Assume that the current flowing isl15 A, and that the electrons travel through copper wire with cross-sectional area 31.2 mm and length 85.5 cm. What is the current density in the wire? The number density of the conduction electrons in copper is 8.49x10 28 /m. Given this number density and the current density, what is the drift speed of the electrons? How long does it take for an 6-15 electron starting at the battery to reach the starter motor? (Ans: 3.69x 106 Am 2.71x10 4 m/s ,52.5 min
Answer #1

(a) The magnitude of current density is given by

       \left | \vec{J} \right |=\left | nq\vec{v} \right |

The direction of \vec{J}and \vec{v} is same for positive charge carriers and opposite for negative charge carriers.

(b) Given:

    I=115A, A=31.2mm2,   L=85.5cm, n=8.49*1028/m3

   The drift velocity of the electron is given by

    v=\frac{I}{neA}

\Rightarrow v=\frac{115}{8.49*10^{28}*1.6*10^{-19}*31.2*10^{-6}}=2.71*10^{-4}m/s

The current density in the wire is

   J=nqv=8.49*10^{28}*1.6*10^{-19}*2.71*10^{-4}=3.69*10^{6}A/m^{2}

The time taken is

t=\frac{L}{v}=\frac{85.5*10^{-2}}{2.71*10^{-4}}=3.15*10^{3}s=52.58min